OpenStudy (anonymous):

Solve the following system by elimination. 3x - 2y = 18 5x + 2y = 14 A. (0, 7) B. (4, -3) C. (6, -6) D. (2, 2)

OpenStudy (anonymous):

whats the answer

OpenStudy (ajspeller):

Add the 2 equations together. That will eliminate the y variables. Then solve for x. Then take the result of x and substitute that back into either of the equations to get the y value. OR Use "brute force" and sub in each x and y value given in the choices. They must result in a true statement for both equations for it to be a valid answer.

OpenStudy (anonymous):

not helping

OpenStudy (dtan5457):

3x-2y =18 5x+2y=14 to solve by elimination, multiply each fraction so that the co efficient of either x or y will cancel out when you add or subtract

OpenStudy (ajspeller):

@dtan5457 that is redundant for this problem. Just add and the y variable goes away.

OpenStudy (dtan5457):

ill even multiply it for you (3x-2y =18)-5 -15x+10y=-90 (5x+2y=14)3 15x+6y=42 add the 2

OpenStudy (dtan5457):

oh wow. did not see that, good eyes, @ajspeller

OpenStudy (ajspeller):

No problem